The ages of the singers of the declaration of Independence are shown below.(Age is approximate since only the birth year appeared in the source, and one has been omitted since his birth year is unknown.) Construct a frequency distribution for the data using seven classes. ( The data for this exercise will be used for exercise 3-23.)

(41,44,44,35,35). (54,52,63,43,46). (47,39,60,48,45). (40,50,27,46,34).
(39,40,42,31,53). (35,30,34,27,50).
(50,34,50,55,50). (37,69,42,63).
(49,39,52,46). (42,45,38,33).
(70,33,36,60). (32,42,45,62).
1
Expert's answer
2020-01-20T09:41:50-0500

Minimum Age = 27

Maximum Age = 70

Range = 70 - 27 = 43

Number of classes = 7

Class width = 43/7 = 6.14 rounded up to next integer as 7.

Sorted data:

27

27

30

31

32

33

33

34

34

34

35

35

35

36

37

38

39

39

39

40

40

41

42

42

42

42

43

44

44

45

45

45

46

46

46

47

48

49

50

50

50

50

50

52

52

53

54

55

60

60

62

63

63

69

70

Frequency distribution: